OpenStudy (anonymous):

i have any question,6(12 ^2+5^2)^0 can you help this question?

OpenStudy (anonymous):

answer is 6 anything to the power 0 is 1 so 6*1 = 6

OpenStudy (anonymous):

how about this question (4^2+3^2)^-1

OpenStudy (anonymous):

Any number [except 0] which is raised to the power 0 is equal to 1 Therefore 6 x 1 = 6 is the result.

OpenStudy (anonymous):

1/25

OpenStudy (anonymous):

(4^2+3^2)^-1 =[16 +9]^-1 = 25^-1 0 1/25

OpenStudy (anonymous):

how about this question 5\p^6

OpenStudy (anonymous):

5p^-6

OpenStudy (anonymous):

how about this question 5\(a+b)^3

OpenStudy (anonymous):

You can only expand the denominator, but I do not see the point in that.

OpenStudy (anonymous):

anyway it is 5/a^3 +a^2b +ab^2 +b^3

OpenStudy (anonymous):

sorry 5/a^3 +2a^2b +2ab^2 +b^3

OpenStudy (anonymous):

sorry,the question is 5/(a+b)^3

OpenStudy (anonymous):

so it is the answer?